Answer Questians 15, 16, and 17 on the basis following information. of the Suppose 50 units of product X can be produced by employing just laborand capital accoiding to thefour tech- niques (A, B, C, and D) shown below. Assume the prices of labor and capital are $5 and $4, respectively. 4) 3 2 Labor 2 3 5 15. Which technique is economically most efficient in pro- ducing product X? (a) A (b) B (c) C (d) D
Answer #1

Option (b).

Total cost (TC) ($) = Labor cost + Capital cost = 5L + 4K

TC, Technique A ($) = 5 x 1 + 4 x 5 = 5 + 20 = 25

TC, Technique B ($) = 5 x 2 + 4 x 3 = 10 + 12 = 22

TC, Technique C ($) = 5 x 3 + 4 x 2 = 15 + 8 = 23

TC, Technique D ($) = 5 x 4 + 4 x 1 = 20 + 4 = 24

Since TC is lowest for technique B, this is the most efficient one.